# Robust Point Cloud Registration via Geometric Overlapping Guided Rotation Search

### v1

@@ Abstract
Point cloud registration based on correspondences computes the rigid transformation that maximizes the number of inliers constrained within the noise threshold. Current state-of-the-art (SOTA) methods employing spatial compatibility graphs or branch-and-bound (BnB) search mainly focus on registration under high outlier ratios. However, graph-based methods require at least quadratic space and time complexity for graph construction, while multi-stage BnB search methods often suffer from inaccuracy due to local optima between decomposed stages. This paper proposes a geometric maximum overlapping registration framework via rotation-only BnB search. The rigid transformation is decomposed using Chasles’ theorem into a translation along rotation axis and a 2D rigid transformation. The optimal rotation axis and angle are searched via BnB, with residual parameters formulated as range maximum query (RMQ) problems. Firstly, the top-k candidate rotation axes are searched within a hemisphere parameterized by cube mapping, and the translation along each axis is estimated through interval stabbing of the correspondences projected onto that axis. Secondly, the 2D registration is relaxed to 1D rotation angle search with 2D RMQ of geometric overlapping for axis-aligned rectangles, which is solved deterministically in polynomial time using sweep line algorithm with segment tree. Experimental results on 3DMatch, 3DLoMatch, and KITTI datasets demonstrate superior accuracy and efficiency over SOTA methods, while the time complexity is polynomial and the space complexity increases linearly with the number of points, even in the worst case.
Index Terms:
Point cloud registration, branch-and-bound (BnB) search, range maximum query (RMQ), interval stabbing, segment tree.
@@ I Introduction
Point cloud registration estimates the rigid transformation with 6 degrees of freedom (DoF) between two point clouds, which is a fundamental problem in simultaneous localization and mapping (SLAM)[1, 2, 3], computer assisted surgery [4, 5, 6] and augmented reality[7, 8, 9]. Although iterative closest point (ICP) [10] and its variants [11, 12, 13] are widely used for fine registration with given initial alignment, they require robust initialization by employing a global registration method.
@@
Global registration methods estimate the rigid transformation between point clouds by establishing correspondences without initial guesses. Conventional correspondence-based approaches rely on handcrafted geometric primitives, such as 4-points congruent sets [14, 15], convex hull [16, 17] and dense local feature descriptors [18, 19]. While generating plausible correspondences, these methods typically produce significant outlier proportions that necessitate robust filtering. Recently, variants of RANSAC [20, 21], graph-based match [22, 23, 24], and branch-and-bound (BnB) search [25, 26, 27, 28] have demonstrated effectiveness in robust global registration against outliers. The exceptional outlier rejection capability of spatial compatibility (SC) from graph-based methods has prompted their integration into other pipelines [21, 29]. Nevertheless, these methods face a critical trade-off: no universal approach simultaneously achieves robustness across varying outlier ratios and correspondence counts while guaranteeing polynomial time and linear space complexity.
@@
To achieve globally optimal point cloud registration with guaranteed space and time complexity, we propose a Geometric Maximum Overlapping Registration (GMOR) framework using two-stage rotation-only BnB search. The method employs Chasles’ decomposition theorem [30] to decouple the rigid transformation, as shown in Fig. 1. The rotation axis is searched by cube mapping parameterization in the first stage and the translation along the axis is estimated by 1D range maximum query (RMQ) of overlapped correspondences projected onto the axis. This 1D RMQ adapts the interval stabbing algorithm used in previous works [26, 27, 28, 29, 31]. The second stage searches rotation angles in 1D space, where we relax rotation centers determined by angles to axis-aligned bounding boxes. Maximum inlier estimation is formulated as a 2D RMQ problem over pairwise intersecting rectangles. We reformulate this geometric maximum overlapping search as the stabbing counting query of Klee’s rectangle problems [32, 33], solving it efficiently via sweep line algorithm inspired by Bentley-Ottmann algorithm [34] augmented with segment tree. By combining BnB over Special Orthogonal Group ($\mathrm{SO}(3)$) with these computational geometry-based techniques, our approach achieves polynomial time complexity through dimension reduction while maintaining linear space complexity. The contributions of our work are as follows:
  • A novel two-stage rotation-only BnB framework using Chasles’ decomposition guarantees global convergence in polynomial time with linear space complexity, eliminating graph construction overhead.
@@
  • Cube-mapped hemispherical parameterization for rotation axis search is coupled with adaptive interval stabbing of translation along the axis, progressively tightening intervals through correspondence projections to accelerate RMQ resolution.
@@
  • A deterministic polynomial-time solution for 2D rigid registration is derived by reformulating as 1D rotation angle search with Klee’s rectangle counting, which utilizes sweep line algorithm with segment tree in $O(N\log N)$ time per iteration.
@@ II-A Feature extraction and learning-based registration
A common approach for global registration is extracting the corresponding features between two point clouds to estimate the transformation between them. Rusu et al. [18] propose a Fast Point Feature Histograms (FPFH) descriptor for global registration, which is a handcrafted feature capturing the local geometry of each point invariant to scene context. Learning-based feature descriptors have been widely used for specific scenes, and 3DMatch proposed by Zeng et al. [35] is a pioneering work which provides the first benchmark dataset for learned feature representations. Choy et al. [36] propose Fully Convolutional Geometric Features (FCGF), which is trained separately on indoor/outdoor datasets and achieves higher inlier ratio. Huang et al. [37] demonstrate that using pairwise point clouds as joint input extracts richer features, proposing Predator for low-overlap cases. Qin et al. [38] further employ Transformer networks to model global geometric relationships of pairwise point clouds.
Based on the correspondence matching in feature space, the rigid transformation is estimated by either RANSAC [39] or chunk-based methods [38]. Recently, end-to-end learning approaches, e.g. DGR [40] and PointDSC [22], are progressive for solving the rigid transformation of correspondences. While effective in specific scenarios [41, 42], they generalize poorly to practical applications. Crucially, traditional methods [21, 43, 24, 44, 29] without learning now surpass them in accuracy and efficiency through graph or solution space exploration. Given these limitations, this section focuses on heuristic global registration techniques.
@@ II-B Iterative Random Sampling
RANSAC [39] is a widely used algorithm for robust estimation with outliers, has inspired numerous variants [45, 20, 21]. Rusu et al. [18] leverage pairwise point distances and Fast Point Feature Histograms (FPFH) descriptor similarities to formulate SAmple Consensus Initial Alignment (SAC-IA), which suppresses outlier influence through finite iterations. Barath and Matas [45] improve sampling efficiency with Graph-Cut RANSAC (GC-RANSAC), removing outliers by graph cutting with local optimization. Unlike local sampling approaches, Li et al. [20] develop a one-point RANSAC variant using global pairwise distance and a modified Tukey’s biweight function for scaled transformation estimation. Shi et al. [21] extend this framework and propose TCF, implementing a three-stage RANSAC process for robust outlier rejection.
@@ II-C Spatial Compatibility Graph Matching
SC [22, 23, 24] is a practical measure of the geometric consistency between point clouds after rigid transformation. SC graph is constructed by pair-wise distances between correspondences, and the vertices are connected if the distances difference is less than a threshold. Yang et al. [23] propose SAC-COT by randomly sampling three-point correspondences guided by the SC graph, which generates more reliable correspondences than RANSAC [39]. Based on reliable seeds selection of SC graph in PointDSC [22], Chen et al. [24] eliminate the training procedure and propose SC2-PCR with a second order graph (SOG) to sample the correspondences, later extended to SC2-PCR++ [46] via consensus reselection with one-to-many matching. These works demonstrate traditional SOG matching’s superiority over end-to-end approaches.
For outlier pruning, Yang et al. [43] propose TEASER++ that combines truncated least squares estimation with practical maximum clique algorithm [47]. Zhang et al. [44] relax the constraint of the maximum clique in TEASER++ [43] and propose MAC to find the maximal cliques of the SOG. This approach employs modified Bron-Kerbosch algorithm implemented in igraph library [48] with complexity $O(d_{v}(N\!-\!d_{v})3^{d_{v}/3})$, where $d_{v}$ is the graph degeneracy. While efficient for sparse graphs, maximal clique-based methods approach exponential space and time complexity on recursively searching in dense graphs, which leads to memory limitations reported in [28].
@@ II-D BnB Search in Solution Space
BnB search provides globally optimal solutions for combinatorial optimization problems, but the time complexity increases exponentially in the dimensionality of the solution space. Recent works [25, 26, 27, 28] focus on the search in the 6-DoF solution space of rigid transformation. Hartley and Kahl [49] pioneered this approach, formulating rotation search as BnB over the closed ball with radius $\pi$ in $\mathrm{R}^{3}$ using Rodrigues parameterization. Yang et al. [25] propose Go-ICP with nested BnB search of $\mathrm{SE}(3)$ to estimate the 6-DoF rigid transformation between two point clouds. To reduce the dimensions of search space, Bustos et al. [50] propose a fast rotation search with circular R-tree indexing, but it still requires the nested $\mathbb{R}^{3}$ translation search. In further research, Bustos and Chin [26] propose GORE, combining BnB with 1D interval stabbing for outlier removal while leaving the nested translation search problem unresolved.
@@
In this paper, our approach is a variant of BnB search that mitigates local optima by top-$k$ rotation axes selection. Compared with the decomposed stages in TR-DE [27], we implement the following key innovations: In stage I, the rotation axis is discretized via spherical cube mapping, and the translation is solved using tighter boundaries with interval stabbing. In stage II, a novel deterministic 2D registration through sweep line algorithm with segment tree compresses the 3-DoF search in TR-DE [27] to 1 DoF, while maintaining registration accuracy.
@@ III Problem Formulation
where $\mathbb{I}(\cdot)$ is the indicator function of inliers, and $w_{i}$ represents the $i^{th}$ correspondence weight in $\mathbf{w}=\{w_{i}\}_{i=1}^{N}$. When $\mathbf{w}=\{1\}^{N}$, the reduces to the maximizing the inliers count. $\xi$ is the noise threshold, following $\xi^{2}=2\sigma^{2}\chi_{3,0.95}^{2}$ from Gaussian noise $\mathcal{N}(0,\sigma^{2}\mathbf{I})$ per dimension of a single point, with $\chi_{3,0.95}^{2}$ being the $95\%$ quantile of the 3-DoF chi-square distribution.
Figure 1: Decomposition of rigid transformation by Chasles’ theorem. (a) Equivalent screw motion from $\mathbf{P}$ to $\mathbf{Q}$; (b) Decomposed translation along the rotation axis direction $\mathbf{r}$; (c) Residual 2D rotation about the axis through a determined center $\mathbf{C}$.
As shown in Fig. 1, the rigid transformation is equivalent to a screw motion, which can be decomposed to a translation along the unit rotation axis direction $\mathbf{r}$ and a rotation about the axis through a fixed point $\mathbf{C}$ satisfying $\mathbf{r}\times(\mathbf{R}\mathbf{C}+\mathbf{t}-\mathbf{C})=\mathbf{0}$. Applying the Rodrigues formula, we reformulate (1) as:
@@
$\mathbf{R}$ $=\mathbf{I}\cos\theta+(1-\cos\theta)\mathbf{r}\mathbf{r}^{T}+\mathbf{r}_{\times}\sin\theta,$ (3b)
@@
$\mathbf{C}$ $=\frac{1}{2}\left(\mathbf{I}-\mathbf{r}\mathbf{r}^{T}+\mathbf{r}_{\times}\cot\left(\frac{\theta}{2}\right)\right)\mathbf{t},$ (3d)
where $\mathbf{S}_{i}$ is the difference vector between $\mathbf{Q}_{i}$ and $\mathbf{P}_{i}$, $\mathbf{I}$ is the identity matrix, $\mathbf{r}_{\times}$ is the skew-symmetric matrix of unit rotation axis $\mathbf{r}$, and $\theta$ is the rotation angle in the Rodriguez formula (3b). Particularly, $\mathbf{C}$ represents an arbitrary point on the rotation axis with direction $\mathbf{r}$, and the selected $\mathbf{C}$ in (3d) is the origin point projected onto the axis. It is easy to prove that (2) is equivalent to (1).

### v2

# GMOR: A Lightweight Robust Point Cloud Registration Framework via Geometric Maximum Overlapping

@@ Abstract
Point cloud registration based on correspondences computes the rigid transformation that maximizes the number of inliers constrained within the noise threshold. Current state-of-the-art (SOTA) methods employing spatial compatibility graphs or branch-and-bound (BnB) search mainly focus on registration under high outlier ratios. However, graph-based methods require at least quadratic space and time complexity for graph construction, while multi-stage BnB search methods often suffer from inaccuracy due to local optima between decomposed stages. This paper proposes a geometric maximum overlapping registration framework via rotation-only BnB search. The rigid transformation is decomposed using Chasles’ theorem into a translation along rotation axis and a 2D rigid transformation. The optimal rotation axis and angle are searched via BnB, with residual parameters formulated as range maximum query (RMQ) problems. Firstly, the top-k candidate rotation axes are searched within a hemisphere parameterized by cube mapping, and the translation along each axis is estimated through interval stabbing of the correspondences projected onto that axis. Secondly, the 2D registration is relaxed to 1D rotation angle search with 2D RMQ of geometric overlapping for axis-aligned rectangles, which is solved deterministically in polynomial time using sweep line algorithm with segment tree. Experimental results on indoor 3DMatch/3DLoMatch scanning and outdoor KITTI LiDAR datasets demonstrate superior accuracy and efficiency over SOTA methods, while the time complexity is polynomial and the space complexity increases linearly with the number of points, even in the worst case.
@@ I Introduction
Point cloud registration estimates the rigid transformation with 6 degrees of freedom (DoF) between two point clouds, which is a fundamental problem in simultaneous localization and mapping (SLAM) [3, 54, 66, 46], computer assisted surgery [37, 82] and augmented reality-based navigation [57, 52, 64, 74]. Although iterative closest point (ICP) [8] and its variants [50, 34, 63, 62] are widely used for fine registration with given initial alignment, they require robust initialization by employing a global registration method.
@@
Global registration methods estimate the rigid transformation between point clouds by establishing correspondences without initial guesses. Conventional correspondence-based approaches rely on handcrafted geometric primitives, such as 4-points congruent sets [1, 45], convex hull [22, 23] and dense local feature descriptors [55, 59]. Although these methods can generate plausible correspondence sets, they usually suffer from a high outlier ratio, which requires robust filtering. To address this challenge, various robust global registration methods have been developed. Random sample consensus (RANSAC) [24] and its variants [36, 58] remain widely used due to their simplicity, while graph-based matching methods [4, 72, 15] apply spatial compatibility (SC) constraints to achieve strong outlier rejection. Deterministic approaches based on branch-and-bound (BnB) search [71, 49] further provide global optimality guarantees in the searched solution space while suffer from rapidly growing search complexity in high-dimensional transformation space.
Within this context, recent deterministic BnB methods reduce the search complexity of rigid registration by decomposing the original 6 DoF into lower-dimensional subproblems [13, 38, 32, 31]. For example, TR-DE [13] adopts a two-stage pure BnB strategy with 3+3 DoF decomposition. Meanwhile, interval stabbing algorithms, which originally designed to maximize a uniform inlier count, are commonly used as 1-DoF reductions within the stages [49, 38, 32, 31]. Chasles’ decomposition theorem [12] offers another practical way to shrink the search space and was first applied to gravity-constrained 4-DoF registration by Li et al. [40]. This method considers a 1+2+1 DoF decomposition and solves the resulting subproblems by interval stabbing, 2D rotation center BnB search, and 1D rotation angle voting. Despite their effectiveness, these multi-stage methods introduce two limitations: the uniform inlier count treats all correspondences equally and do not incorporate correspondence weights, which reduces discrimination when correspondence confidence is highly non-uniform; and sequentially decomposed stages may propagate local optima between stages. Moreover, since interval stabbing collapses only a single DoF, it remains challenging to design a unified reduction that systematically guides BnB toward further dimension reduction.
@@
To mitigate these limitations while maintaining efficient runtime and low memory consumption, we propose a Geometric Maximum Overlapping Registration (GMOR) framework based on a two-stage rotation-only BnB search. The method leverages Chasles’ decomposition theorem [12, 33] to decouple rigid motion in the general 6-DoF case, as illustrated in Fig. 1. Although the resulting DoF decomposition is similar in spirit to TR-DE [13] and the gravity-constrained formulation in [40], our framework differs in its rotation-only search space and algorithmic structure. Constrained settings, such as gravity-constrained 4-DoF registration [10, 40], naturally arise as special cases within this general framework. In the first stage, the rotation axis is searched directly on the unit sphere using a cube-mapping hemispherical parameterization, and the translation along the axis is estimated via a 1D range maximum query (RMQ) over projected correspondences. This RMQ adapts interval stabbing algorithms from prior works [49, 32, 31, 77, 38, 40], but is used here to maximize a weighted correspondence overlap objective rather than uniform inlier counting. In the second stage, the residual rotation angle is searched in 1D space and the rotation axes with top-k weights in the first stage are selected to mitigate the local optima between stages. Rotation centers induced by candidate angles are relaxed into axis-aligned bounding boxes, and maximum overlap estimation is formulated as a 2D RMQ problem over pairwise intersecting rectangles. We further reformulate this geometric maximum overlapping as a stabbing counting query of Klee’s rectangle problem [19, 6], which can be solved deterministically in linear-logarithmic time using sweep line algorithm inspired by Bentley-Ottmann algorithm [7] augmented with a segment tree.
This viewpoint provides a unified interpretation across both stages: casting interval stabbing as a solver of 1D RMQ problem also motivates the 2D RMQ formulation in the second stage, further reducing the BnB search dimension. The proposed framework combines rotation-only BnB search over the Special Orthogonal Group ($\mathrm{SO}(3)$) with computational geometry-based RMQ solvers to achieve effective dimension reduction while preserving polynomial time complexity and linear space complexity. The main contributions of this work are summarized as follows:
  • We propose a novel two-stage rotation-only BnB framework for general 6-DoF rigid registration, which optimizes a weighted correspondence overlap objective rather than uniform inlier counting.
@@
  • A cube-mapping hemispherical parameterization is introduced for global rotation axis search, coupled with adaptive interval stabbing to estimate translation along the axis. The correspondence projections progressively tighten the search bounds, enabling efficient overlap maximization.
@@
  • We derive a deterministic polynomial-time solution for the residual 2D rigid registration by reformulating it as a 1D rotation angle search with 2D RMQ. The resulting geometric overlap maximization is cast as a Klee’s rectangle problem [19] and solved via a sweep line algorithm with segment tree in $O(N\log N)$ time per iteration.
@@ II-A Feature extraction and learning-based registration
A common approach for global registration is extracting the corresponding features between two point clouds to estimate the transformation between them. Rusu et al. [55] propose a Fast Point Feature Histograms (FPFH) descriptor for global registration, which is a handcrafted feature capturing the local geometry of each point invariant to scene context. Learning-based feature descriptors have been widely used for specific scenes, and 3DMatch proposed by Zeng et al. [76] is a pioneering work which provides the first benchmark dataset for learned feature representations. Choy et al. [17] propose Fully Convolutional Geometric Features (FCGF), which is trained separately on indoor/outdoor datasets and achieves higher inlier ratio. Huang et al. [30] demonstrate that using pairwise point clouds as joint input extracts richer features, proposing Predator for low-overlap cases. Qin et al. [53] further employ Transformer networks to model global geometric relationships of pairwise point clouds. To address the positional deviation problem in predicted correspondences, Zhao et al. [79] propose a dual-branch network for matching and predicting patch-based features respectively.
Based on the correspondence matching in feature space, the rigid transformation is estimated by either RANSAC [24] or chunk-based methods [53]. Recently, end-to-end learning approaches, e.g. DGR [16] and PointDSC [4], are progressive for solving the rigid transformation of correspondences. While effective in specific scenarios [65, 73], they generalize poorly to practical applications. Crucially, traditional methods [58, 70, 15, 78, 31] without learning now surpass them in accuracy and efficiency through graph or solution space exploration. These observations motivate a trend towards front-end and back-end decoupling [51, 11, 43], where the front end focuses on feature descriptor and correspondence generation, while the back end is responsible for reliable robust estimation. Following this paradigm, this section concentrates on back-end global registration techniques.
@@ II-B Iterative Random Sampling
RANSAC [24] is a widely used algorithm for robust estimation with outliers, has inspired numerous variants [5, 36, 58]. Rusu et al. [55] leverage pairwise point distances and Fast Point Feature Histograms (FPFH) descriptor similarities to formulate SAmple Consensus Initial Alignment (SAC-IA), which suppresses outlier influence through finite iterations. Barath and Matas [5] improve sampling efficiency with Graph-Cut RANSAC, removing outliers by graph cutting with local optimization. Unlike local sampling approaches, Li et al. [36] develop a one-point RANSAC variant using global pairwise distance and a modified Tukey’s biweight function for scaled transformation estimation. Shi et al. [58] extend this framework and propose TCF, implementing a three-stage RANSAC process for robust outlier rejection.
@@ II-C Spatial Compatibility Graph Matching
SC [4, 72, 15] is a practical measure of the geometric consistency between point clouds after rigid transformation. SC graph is constructed by pair-wise distances between correspondences, and the vertices are connected if the distances difference is less than a threshold. Yang et al. [72] propose SAC-COT by randomly sampling three-point correspondences guided by the SC graph, which generates more reliable correspondences than RANSAC [24]. Yan et al. [68] optimize the candidate selection by the reliability degree of graph nodes and edges for outlier removal. Based on reliable seeds selection of SC graph in PointDSC [4], Chen et al. [15] eliminate the training procedure and propose SC2-PCR with a second order graph (SOG) to sample the correspondences, later extended to SC2-PCR++ [14] via consensus reselection with one-to-many matching and mixed feature-spatial consistency metric. Yan et al. [69] consider the GPU-based parallel searching and propose pivot-guided search algorithm for SOG. These works demonstrate traditional SOG matching’s superiority over end-to-end approaches.
For outlier pruning, Yang et al. [70] propose TEASER++ that combines truncated least squares estimation with practical maximum clique algorithm [47]. Zhang et al. [78] relax the constraint of the maximum clique in TEASER++ [70] and propose MAC to find the maximal cliques of the SOG. This approach employs modified Bron-Kerbosch algorithm implemented in igraph library [21] with complexity $O(d_{v}(N\!-\!d_{v})3^{d_{v}/3})$, where $d_{v}$ is the graph degeneracy. Laserna [35] proposes CliReg for sparse graphs, which adopts BnB algorithm in maximal clique search of sparse graph. While efficient for sparse graphs, maximal clique-based methods approach exponential space and time complexity on recursively searching in dense graphs, which leads to memory limitations reported in [32].
@@ II-D BnB Search in Solution Space
BnB search provides globally optimal solutions for combinatorial optimization problems, but the time complexity increases exponentially in the dimensionality of the solution space. Recent works [71, 49, 13, 32] focus on the search in the 6-DoF solution space of rigid transformation. Hartley and Kahl [27] pioneered this approach, formulating rotation search as BnB over the closed ball with radius $\pi$ in $\mathrm{R}^{3}$ using Rodrigues’ parameterization. Yang et al. [71] propose Go-ICP with nested BnB search of $\mathrm{SE}(3)$ to estimate the 6-DoF rigid transformation between two point clouds. To reduce the dimensions of search space, Bustos et al. [48] propose a fast rotation search with circular R-tree indexing, but it still requires the nested $\mathbb{R}^{3}$ translation search. In further research, Bustos and Chin [49] propose GORE, combining BnB with 1D interval stabbing for outlier removal while leaving the nested translation search problem unresolved.
@@
II-E Practical LiDAR registration with Gravity Priors
In contrast to the general 6-DoF registration problem, several studies [10, 39, 40, 41] have investigated a practical 4-DoF setting enabled by gravity priors in LiDAR systems, where the rotation is constrained to the yaw angle. With roll and pitch approximately aligned with the gravity direction, the solution space is effectively reduced to a 2.5D problem, consisting of 3 translational and 1 rotational DoF.
Cai et al. [10] study this setting by adopting Intrinsic Shape Signatures (ISS) [80] for correspondence generation and propose Fast Match Pruning (FMP) to further suppress outliers. The resulting transformation is estimated via 3-DoF BnB search over the translation space with 1-DoF rotation angle interval stabbing. Li et al. [39] also perform the translation space BnB search but estimate the remaining rotation angle through global voting. To further improve efficiency based on [39], Li et al. [40] decompose the 4-DoF problem into multiple stages using Chasles’ theorem [12]. Specifically, the translation along the known rotation axis is first estimated via interval stabbing, followed by a 2D BnB search for the rotation center, and a final global voting step to recover the rotation angle. Lim et al. [41] consider the 4-DoF case of TEASER++ [70] and propose Quatro, which adopts maximal clique in SC and graduated non-convexity based truncated least square estimation of Quasi-$\mathrm{SO}(3)$ solution. Recently, Aoki et al. [2] propose a correspondence-free formulation that performs global BnB search jointly over translation and yaw angle, assuming roll and pitch lie within a small neighborhood around zero. This approach avoids explicit correspondence matching in feature space but relies on the voxel map-based nearest neighbor search at the cost of increased computational complexity. Nevertheless, the above BnB-based methods treat translation as the search domain, making their search efficiency and resolution sensitive to the spatial scale of the point clouds.
Building upon the 2.5D formulation, an even more aggressive dimensionality reduction has been adopted by representing LiDAR point clouds in a 2D bird’s-eye-view (BEV) space [67, 75, 44], which enables efficient point cloud retrieval for loop detection. By projecting point clouds onto the ground plane and discarding the dimension along gravity direction, the registration problem is further simplified to a purely 2D formulation with 3 DoF, significantly reducing computational complexity while retaining sufficient geometric structure for localization and alignment [29]. Overall, the 4-DoF and 3-DoF registration problems can be viewed as special cases of the general 6-DoF formulation, which can be flexibly addressed within a unified BnB framework by appropriately redefining the search domain.
@@ III Problem Formulation
where $\mathbb{I}(\cdot)$ is the indicator function of inliers, and $w_{i}$ represents the $i^{th}$ correspondence weight in $\mathbf{w}=\{w_{i}\}_{i=1}^{N}$. When $\mathbf{w}=\{1\}^{N}$, the reduces to the maximizing the inliers count. $\xi$ is the noise threshold, following $\xi^{2}=2\sigma^{2}\chi_{3,0.95}^{2}$ from Gaussian noise $\mathcal{N}(0,\sigma^{2}\mathbf{I})$ per dimension of a single point, with $\chi_{3,0.95}\approx 2.796$ being the $95\%$ quantile of the 3-DoF chi-square distribution.
Figure 1: Decomposition of rigid transformation by Chasles’ theorem [12]. (a) Equivalent screw motion from $\mathbf{P}$ to $\mathbf{Q}$; (b) Decomposed translation along the rotation axis direction $\mathbf{r}$; (c) Residual 2D rotation about the axis through a determined center $\mathbf{C}$.
As shown in Fig. 1, the rigid transformation is equivalent to a screw motion, which can be decomposed to a translation along the unit rotation axis direction $\mathbf{r}$ and a rotation about the axis through a fixed point $\mathbf{C}$ satisfying $\mathbf{r}\times(\mathbf{R}\mathbf{C}+\mathbf{t}-\mathbf{C})=\mathbf{0}$. Applying the Rodrigues’ formula, we reformulate (1) as:
@@
$\mathbf{R}$ $=\mathbf{I}\cos\theta+(1-\cos\theta)\mathbf{r}\mathbf{r}^{\top}+\mathbf{r}_{\times}\sin\theta,$ (3b)
@@
$\mathbf{C}$ $=\frac{1}{2}\left(\mathbf{I}-\mathbf{r}\mathbf{r}^{\top}+\mathbf{r}_{\times}\cot\left(\frac{\theta}{2}\right)\right)\mathbf{t},$ (3d)
where $\mathbf{S}_{i}$ is the difference vector between $\mathbf{Q}_{i}$ and $\mathbf{P}_{i}$, $\mathbf{I}$ is the identity matrix, $\mathbf{r}_{\times}$ is the skew-symmetric matrix of unit rotation axis $\mathbf{r}$, and $\theta$ is the rotation angle in the Rodrigues’ formula (3b). Particularly, $\mathbf{C}$ represents an arbitrary point on the rotation axis with direction $\mathbf{r}$, and the selected $\mathbf{C}$ in (3d) is the origin point projected onto the axis. The derivation from (1) to (2) is presented in Appendix A.